OpenStudy (anonymous):

Evaluate sixteen to the three-halves power 64 4 12 9

OpenStudy (slaaibak):

\[16^{3\over2} = ({2^4})^{3/2} = 2^6 = 64\]

OpenStudy (anonymous):

Thanks for showing the steps! very helpful :)